OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = x3 + 4 and g(x) = 3√(x-4)

OpenStudy (anonymous):

It is simple. Give it a try. If it doesn't work out, we'll do it.

OpenStudy (marissalovescats):

What you do is plug each equation into each others x. So to do: f(g(x)) You plug the equation of g(x) into the x place of f(x) and it should simplify down to x

OpenStudy (anonymous):

y=x3+4 y-4=x3 3√(y-4)=x?

OpenStudy (marissalovescats):

 See how I put the function of g(x) into the x place of f(x)?